# MotionStream: Real-Time Video Generation with Interactive Motion Controls

### v1

@@ 1 Introduction
The ultimate goal of motion-controlled video synthesis is to give creators the power of a director’s chair, allowing them to intuitively guide digital actors, objects, and cameras in real time. Although recent video diffusion models have made impressive strides toward this goal (Wang et al., 2024b; Geng et al., 2025; Burgert et al., 2025; Zhang et al., 2025b; Li et al., 2025b; 2024a; Niu et al., 2024; Shi et al., 2024a; Zhou et al., 2025a; Lei et al., 2025), generating high-fidelity videos following user-specified motion trajectories, the experience remains far from interactive.
The promise of interactive control is currently hindered by several fundamental constraints. First, generation is too slow for interaction. For example, synthesizing a 5-second video clip with Motion Prompting (Geng et al., 2025) takes 12 minutes, trapping users in frustrating “render-and-wait” cycles. Second, the process is inherently non-causal, since diffusion models process the entire sequence in parallel with bidirectional attention. A user cannot see any partial results until the entire motion specification is complete. Finally, the inability to generate more than a few seconds of video severely limits the scope for any meaningful or extended creative expression. Together, these constraints (slow, non-causal, and short-duration generation) undermine the potential for a truly interactive creative experience.
@@
Our approach starts with a motion-controlled teacher model that uses lightweight sinusoidal embeddings with channel-wise concatenation for trajectory conditioning, avoiding the computational overhead of ControlNet-style (Zhang et al., 2023) architectures. Trained on both text and motion conditions, we introduce joint text-motion guidance that balances precise trajectory adherence with natural secondary motions enabled by text prompts. We then distill this teacher into a causal student through Self Forcing-style self-rollout (Huang et al., 2025b). While effective for short sequences, standard approaches drift during extended generation. Our analysis of attention patterns reveals persistent focus on initial frames alongside local temporal dependencies, similar to StreamingLLM (Xiao et al., 2023). This insight drives our attention sinking mechanism with rolling KV caches, which we incorporate directly into training to properly simulate inference-time extrapolation distributions, ensuring stable, indefinite-length generation at constant latency through fixed context windows.
@@ 2 Related Work
Controllable Video Generation. Enabling precise user control is essential for applying video generative models to diverse downstream applications (Li et al., 2025c; Tu et al., 2025; Bahmani et al., 2024a; Gao et al., 2024c; Wu et al., 2025b; Fu et al., 2025). To this end, a large body of recent research has explored various types of control signals for video generation, such as structure control (Xing et al., 2024; Yang et al., 2025a; Jiang et al., 2025; Pang et al., 2024; Xing et al., 2025), camera control (Gao et al., 2024b; Zheng et al., 2024; He et al., 2024; Bai et al., 2025; Wu et al., 2025a; YU et al., 2025; Bahmani et al., 2024b; Yang et al., 2024b; Zheng et al., 2024; 2025), subject control (Huang et al., 2025a; Liu et al., 2025b; Fei et al., 2025; Liu et al., 2025a), and audio control (Tian et al., 2024; Gao et al., 2025; Peng et al., 2024).
@@
As a unique modality that captures underlying video dynamics, motion has become a key conditioning signal for recent video diffusion models. Recent video diffusion models often condition generated videos on diverse forms of motion representations, including optical flow, 2D/3D motion trajectories, bounding boxes, and semantic segmentation (Goldman et al., 2008; Niu et al., 2024; Li et al., 2024b; Wu et al., 2024b; Geng et al., 2025; Zhang et al., 2025b; Gillman et al., 2025; Shi et al., 2024a; Wu et al., 2024a; Gu et al., 2025b; Burgert et al., 2025; Tanveer et al., 2024). Despite their impressive quality, these methods are fundamentally limited to offline processing because they rely on diffusion models with full bidirectional attention, which requires the entire control signal to be known in advance. This constraint prevents their use in real-time, interactive applications.
@@
Autoregressive Video Models. Early work adopted generative adversarial networks (GANs) for autoregressive or parallel video synthesis (Vondrick et al., 2016; Brooks et al., 2022; Villegas et al., 2017; Denton et al., 2017; Tulyakov et al., 2018; Liu et al., 2021; Li et al., 2022). More recently, there has been a paradigm shift towards using diffusion models trained with denoising objectives (Ho et al., 2022; Blattmann et al., 2023b; Yang et al., 2025b; Kong et al., 2024; Polyak et al., 2024; Blattmann et al., 2023a; Villegas et al., 2023; Deng et al., 2025; Gupta et al., 2024; Wang et al., 2025a), or autoregressive (AR) models trained with next-token prediction (Weissenborn et al., 2020; Kondratyuk et al., 2024; Yan et al., 2021; Wang et al., 2024a; Bruce et al., 2024; Ren et al., 2025).
@@
Another line of research integrates AR and diffusion to enable causal, high-quality video generation (Gu et al., 2025a; Ruhe et al., 2024; Kim et al., 2024; Xie et al., 2024; Zhang & Agrawala, 2025; Sun et al., 2025; Weng et al., 2024; Liu et al., 2024; Chen et al., 2024; Guo et al., 2025b; Hu et al., 2024; Jin et al., 2025; Gu et al., 2025a; Gao et al., 2024a; Li et al., 2025e; Zhang et al., 2025a). Our work is inspired by the recent paradigm that distills a slow teacher model into a fast AR student for real-time performance (Yin et al., 2025; Huang et al., 2025b; Lin et al., 2025). However, these approaches either exhibit severe color drifts beyond the training horizon or require complex long-video finetuning, which poses challenges for controllable video generation.
@@
Interactive Video World Model. Our work also belongs to interactive video world models, which aim to simulate environments for real-time interaction. This area has recently gained significant attention, as several recent works have demonstrated impressive real-time, user-driven interaction (Ball et al., 2025; Li et al., 2025a; Team, 2025; He et al., 2025; Bar et al., 2025; Po et al., 2025). However, most existing approaches either require substantial compute for inference (Ball et al., 2025; Parker-Holder et al., 2024), or are limited to closed-domain or synthetic environments (Yu et al., 2025; Guo et al., 2025a; Yang et al., 2024a). In contrast, our work demonstrates that real-time, interactive generation for open-domain, photorealistic videos can be achieved on a single GPU.
@@ 3 MotionStream: Streaming Generation Meets Motion Controls
Existing motion-conditioned video generation methods achieve strong motion-video alignment, but cannot support streaming interaction since bidirectional attention requires all future control signals upfront. Our proposed MotionStream addresses this through carefully designed causal distillation techniques, as illustrated in Figure 2. We first describe how to equip a pretrained video diffusion model with motion-control capability (Sec. 3.1) to serve as our bidirectional teacher model, then introduce our training and inference pipeline for real-time streaming via causal distillation (Sec. 3.2).
Figure 2: Model architecture and training pipeline. To build a teacher motion-controlled video model, we extract and randomly sample 2D tracks from the input video and encode them using a lightweight track head. The resulting track embeddings are combined with the input image, noisy video latents, and text embeddings as input to the diffusion transformer with bidirectional attention, which is then trained with a flow matching loss (top). We then distill a few-step causal diffusion model from the teacher through Self Forcing-style DMD distillation, where autoregressive rollout with rolling KV cache and attention sink is applied during both training and inference (bottom).
@@ 3.1 Adding Motion Controls to Bidirectional Teacher Models
Training a high-quality motion-conditioned teacher model is important, as it determines both the quality upper bound and the architectural efficiency of the final distilled system. The teacher must also support diverse motion modalities, from complex real-world object dynamics to camera motions and user drags, which we achieve through the following design. We build our motion-guided teacher model on top of the Wan DiT family (Wang et al., 2025a).
Track Representation and Track Head Design. Following MotionPrompting (Geng et al., 2025), each 2D track is assigned a unique $d$-dimensional embedding vector $\phi_{n}$, derived from a randomly sampled ID number through sinusoidal positional encoding. While encoding tracks as an RGB video and processing it through the CausalVAE is possible, we find that representing them with sinusoidal embeddings with a learnable track head achieves superior track adherence, video quality, and faster speed. We validate this in our experiments in Table 3. Given $N$ tracks $\{(x_{t}^{n},y_{t}^{n})\}_{n=1}^{N}$ across $T$ temporal frames, the input track-conditioning signal $c_{m}\in\mathbb{R}^{T\times H/s\times W/s\times d}$ is constructed by placing visible track embeddings at spatially downsampled locations, where $s$ is the VAE spatial downsampling rate and $v[t,n]\in\{0,1\}$ indicates track visibility:
@@
Our lightweight track-encoding head performs $4\times$ temporal compression followed by a $1\times 1\times 1$ convolution. Prior methods adopt a ControlNet-style architecture (Zhang et al., 2023; Geng et al., 2025), which doubles FLOPs by duplicating network blocks. Instead, we directly concatenate the processed track embeddings with video latents, requiring only minor channel adjustments in the patchifying layer while leaving the core DiT architecture unchanged.
Training. We train the motion-guided teacher model through rectified flow matching objective (Liu et al., 2022; Lipman et al., 2022), where the forward process linearly interpolates between data $z_{0}$ and Gaussian noise $z_{1}\sim\mathcal{N}(0,I)$: $z_{t}=(1-t)z_{0}+tz_{1},\ t\in[0,1]$. The model is trained to predict the expected velocity fields with conditional flow matching loss $\mathcal{L}_{\text{FM}}$. One important limitation to note is that the model cannot inherently distinguish between occluded (non-visible) tracks and unspecified tracks, as both are represented by zero values. When a user releases controls during interaction, the model cannot determine whether the sudden zero values indicate occlusion or simply the absence of specification. This ambiguity occasionally leads to artifacts where objects abruptly appear or disappear. To address this issue, we introduce stochastic mid-frame masking with probability $p_{\text{mask}}=0.2$, where $c[t_{\text{rand}},:,:]=0$, for randomly selected mid-frame chunks $t_{\text{rand}}$. In practice, we first train the model without masking to establish strong track-following capability, and then fine-tune with stochastic masking to preserve coherence when track signals change intermittently.
@@
where $v_{\text{base}}=\alpha\cdot v(\varnothing,c_{m})+(1-\alpha)\cdot v(c_{t},\varnothing)$ and $\alpha=w_{t}/(w_{t}+w_{m})$. We find that the joint guidance weights ($w_{t}=3.0$, $w_{m}=1.5$) provide a good balance: text conditioning enables realistic dynamics, even with sparse, flat-grid inputs, while track guidance preserves trajectories and maintains shape fidelity. We further analyze these effects in Sec. 4.3. Although this increases sampling cost from $2$ to $3$ function evaluations (NFE) per denoising step in the teacher model, our causal distillation (described in the next section) embeds all guidance into a single NFE, thereby eliminating this overhead in the student model.
@@ 3.2 Causal Distillation
Existing motion-controlled video diffusion models require around 50 denoising steps to generate high-quality video and also presume that the entire input motion trajectory is known before generation. In this section, we distill the slow teacher model into a causal video diffusion model, enabling real-time streaming of long videos with motion control. Our training pipeline starts from the Self Forcing paradigm Huang et al. (2025b), but off-the-shelf, we find this exhibits large latency fluctuations due to varying attention window sizes and performs well only within the teacher’s training horizon (i.e., 81 frames), with quality quickly degrading when extrapolating to longer sequences. This section presents several key technical innovations to address these issues.
@@
Attention sink in video diffusion model. Autoregressive generation with sliding-window attention, as employed in Self Forcing (Huang et al., 2025b), is prone to quality degradation and drift during long-video extrapolation. To understand this failure, we visualize the self-attention maps in Figure 3 for both bidirectional and causal attention. Notably, many attention heads focus on the initial tokens corresponding to the input image. This phenomenon mirrors the observations in large language models Xiao et al. (2023), where initial tokens play a crucial role in stable streaming generation. Inspired by this, we adapt the attention sink concept to our video model by maintaining a local attention window while preserving the initial frame’s tokens as a fixed anchor during both training and inference. In our chunk-wise autoregressive setup, this sinking mechanism prevents the model from drifting during long-video extrapolation, even when trained only on short sequences limited by the teacher model.
Causal Adaptation. Following the initialization protocol from CausVid (Yin et al., 2025), we start our student model with the weights of the motion-guided teacher diffusion model and adapt it for causal attention architecture and few-step trajectory, using regression on ODE solution pairs sampled from the teacher. We train the model using attention mask with varying context window size and attention sink size for better generalization.
Self Forcing-Style Distillation. Following Self Forcing (Huang et al., 2025b), we perform temporal autoregressive roll-out with distribution matching distribution (DMD) (Yin et al., 2024b; a). During roll-out, video chunks are generated sequentially, conditioned on previously self-generated outputs rather than ground truth using KV cache.
@@
In our approach, because the KV cache is continuously updated, RoPE values are assigned based on cache position rather than absolute temporal index. We apply KV cache rolling with attention sinks during both training and inference, implementing local attention without explicit attention masking and fully bridging the train–test gap, even in extrapolation scenarios. While TalkingMachines (Low & Wang, 2025) also explores attention sinks, our method differs in two key aspects that enhance long-video stability. First, we eliminate the train-inference mismatch by explicitly simulating the extrapolation process during training, using self-rollout with a rolling KV cache and attention sinks. In contrast, TalkingMachines employs synchronized denoising with causal attention masks, which does not fully replicate the dynamics of autoregressive inference. Second, our training process ensures the teacher model always evaluates continuous video frames. The setup in TalkingMachines introduces a temporal discontinuity between the sink frame and subsequent frames, which can push the input outside the teacher’s pre-training distribution. By maintaining continuity, our teacher provides more robust scores for distillation.
@@
Intuitively, the driving gradient for the generator comes from the score difference between estimated real and fake scores. As we observed superior performance with joint CFG, we compute the real score using this joint score guidance similar to Eq. 2. However, we do not apply any CFG to the fake score (critic model) to keep the score difference meaningful and informative for generator updates.
We update the generator and fake score function (critic) with a ratio of 1:5 (Yin et al., 2024a), allowing the critic to better approximate generated distributions. To manage GPU memory, we adopt Self Forcing’s gradient truncation strategy: randomly sampling denoising step $k$ from $[1,K]$ and using only that step’s output for self-rollout. This maintains computation graphs only for final denoised embeddings in the KV cache, reducing memory while preserving critical gradient flow.

### v2

@@ 1 Introduction
The ultimate goal of motion-controlled video synthesis is to give creators the power of a director’s chair, allowing them to intuitively guide digital actors, objects, and cameras in real time. Although recent video diffusion models have made impressive strides toward this goal (wang2024motionctrl; geng2025motion; burgert2025go; zhang2025tora; li2025magicmotion; li2024puppet; niu2024mofa; shi2024motion; zhou2025trackgo; lei2025animateanything), generating high-fidelity videos following user-specified motion trajectories, the experience remains far from interactive.
The promise of interactive control is currently hindered by several fundamental constraints. First, generation is too slow for interaction. For example, synthesizing a 5-second video clip with Motion Prompting (geng2025motion) takes 12 minutes, trapping users in frustrating “render-and-wait” cycles. Second, the process is inherently non-causal, since diffusion models process the entire sequence in parallel with bidirectional attention. A user cannot see any partial results until the entire motion specification is complete. Finally, the inability to generate more than a few seconds of video severely limits the scope for any meaningful or extended creative expression. Together, these constraints (slow, non-causal, and short-duration generation) undermine the potential for a truly interactive creative experience.
@@
Our approach starts with a motion-controlled teacher model that uses lightweight sinusoidal embeddings with channel-wise concatenation for trajectory conditioning, avoiding the computational overhead of ControlNet-style (zhang2023adding) architectures. Trained on both text and motion conditions, we introduce joint text-motion guidance that balances precise trajectory adherence with natural secondary motions enabled by text prompts. We then distill this teacher into a causal student through Self Forcing-style self-rollout (huang2025self). While effective for short sequences, standard approaches drift during extended generation. Our analysis of attention patterns reveals persistent focus on initial frames alongside local temporal dependencies, similar to StreamingLLM (xiao2023efficient). This insight drives our attention sinking mechanism with rolling KV caches, which we incorporate directly into training to properly simulate inference-time extrapolation distributions, ensuring stable, indefinite-length generation at constant latency through fixed context windows.
@@
Our key contributions are:
  1. 1.
    We present the first streaming motion-conditioned video generation pipeline capable of running at 29.5 FPS on a single H100 GPU, enabling real-time interactive applications.
  2. 2.
    We propose a synergistic system harmonizing efficient architectural designs, including a lightweight track head and conditioning modules, with a distillation process that integrates joint text-motion guidance into the training objective, further accelerated by a Tiny VAE.
  3. 3.
    We introduce a distillation strategy for long video generation that systematically explores attention sinks and local attention with extrapolation-aware training for the first time, effectively preventing drift during long-term streaming.
  4. 4.
    Our approach achieves state-of-the-art results on motion transfer and camera control at orders of magnitude faster speeds, robustly generalizing to diverse interactive use cases.
@@ 2 Related Work
Controllable Video Generation. Enabling precise user control is essential for applying video generative models to diverse downstream applications (li2025diffueraser; tu2025videoanydoor; bahmani20244d; gao2024vista; wu2025video; fu2025learning). To this end, a large body of recent research has explored various types of control signals for video generation, such as structure control (xing2024tooncrafter; yang2025layeranimate; jiang2025vidsketch; pang2024dreamdance; xing2025motioncanvas), camera control (gao2024cat3d; zheng2024cami2v; he2024cameractrl; bai2025recammaster; wu2025cat4d; yu2025trajectorycrafter; bahmani2024vd3d; yang2024direct; zheng2024cami2v; zheng2025vidcraft3), subject control (huang2025videomage; liu2025animateanywhere; fei2025skyreels; liu2025phantom), and audio control (tian2024emo; gao2025wans2vaudiodrivencinematicvideo; peng2024synctalk).
@@
As a unique modality that captures underlying video dynamics, motion has become a key conditioning signal for recent video diffusion models. Recent video diffusion models often condition generated videos on diverse forms of motion representations, including optical flow, 2D/3D motion trajectories, bounding boxes, and semantic segmentation (goldman2008video; niu2024mofa; li2024generative; wu2024draganything; geng2025motion; zhang2025tora; gillman2025force; shi2024motion; wu2024motionbooth; gu2025diffusion; burgert2025go; tanveer2024motionbridge). Despite their impressive quality, these methods are fundamentally limited to offline processing because they rely on diffusion models with full bidirectional attention, which requires the entire control signal to be known in advance. This constraint prevents their use in real-time, interactive applications.
@@
Autoregressive Video Models. Early work adopted generative adversarial networks (GANs) for autoregressive or parallel video synthesis (vondrick2016generating; brooks2022generating; villegas2017decomposing; denton2017unsupervised; tulyakov2018mocogan; liu2021infinite; li2022infinitenature). More recently, there has been a paradigm shift towards using diffusion models trained with denoising objectives (ho2022video; blattmann2023align; yang2025cogvideox; kong2024hunyuanvideo; polyak2024movie; blattmann2023stable; villegas2023phenaki; deng2024nova; gupta2024photorealistic; wang2025wan), or autoregressive (AR) models trained with next-token prediction (weissenborn2020scaling; kondratyuk2024videopoet; yan2021videogpt; wang2024loong; Bruce2024GenieGI; ren2025next).
@@
Another line of research integrates AR and diffusion to enable causal, high-quality video generation (gu2025long; ruhe2024rolling; kim2025fifo; xie2024progressive; zhang2025packing; sun2025ar; weng2024art; liu2024redefining; chen2024diffusion; guo2025long; hu2024acdit; jin2024pyramidal; gu2025long; gao2024ca2; li2025arlon; zhang2025test). Our work is inspired by the recent paradigm that distills a slow teacher model into a fast AR student for real-time performance (yin2025causvid; huang2025self; lin2025diffusion). However, these approaches either exhibit severe color drifts beyond the training horizon or require complex long-video finetuning, which poses challenges for controllable video generation.
@@
Interactive Video World Model. Our work also belongs to interactive video world models, which aim to simulate environments for real-time interaction. This area has recently gained significant attention, as several recent works have demonstrated impressive real-time, user-driven interaction (ball2025genie; li2025hunyuan; team2025yan; he2025matrix; bar2025navigation; po2025long). However, most existing approaches either require substantial compute for inference (ball2025genie; parkerholder2024genie2), or are limited to closed-domain or synthetic environments (yu2025gamefactory; guo2025mineworld; yang2024playable). In contrast, our work demonstrates that real-time, interactive generation for open-domain, photorealistic videos can be achieved on a single GPU.
@@ 3 MotionStream: Streaming Generation Meets Motion Controls
Existing motion-conditioned video generation methods achieve strong motion-video alignment, but cannot support streaming interaction since bidirectional attention requires all future control signals upfront. Our proposed MotionStream addresses this through carefully designed causal distillation techniques, as illustrated in Figure 2. We first describe how to equip a pretrained video diffusion model with motion-control capability (Sec. 3.1) to serve as our bidirectional teacher, utilizing a lightweight track head and control modules designed to minimize architectural overhead. We then introduce our causal distillation pipeline, which performs extrapolation-aware training with attention sinks and local windows for long video generation, while integrating expensive joint text-motion guidance directly into the distillation objective for efficiency. Combined with our Tiny VAE, these joint efforts enable a highly responsive streaming experience.
Figure 2: Model architecture and training pipeline. To build a teacher motion-controlled video model, we extract and randomly sample 2D tracks from the input video and encode them using a lightweight track head. The resulting track embeddings are combined with the input image, noisy video latents, and text embeddings as input to the diffusion transformer with bidirectional attention, which is then trained with a flow matching loss (top). We then distill a few-step causal diffusion model from the teacher through Self Forcing-style DMD distillation, integrating joint text-motion guidance into the objective, where autoregressive rollout with rolling KV cache and attention sink is applied during both training and inference (bottom).
@@ 3.1 Adding Motion Controls to Bidirectional Teacher Models
Training a high-quality motion-conditioned teacher model is important, as it determines both the quality upper bound and the architectural efficiency of the final distilled system. The teacher must also support diverse motion modalities, from complex real-world object dynamics to camera motions and user drags, which we achieve through the following design. We build our motion-guided teacher model on top of the Wan DiT family (wang2025wan).
Track Representation and Track Head Design. Following MotionPrompting (geng2025motion), each 2D track is assigned a unique $d$-dimensional embedding vector $\phi_{n}$, derived from a randomly sampled ID number through sinusoidal positional encoding. While encoding tracks as an RGB video and processing it through the CausalVAE is possible, we find that representing them with sinusoidal embeddings with a learnable track head achieves superior track adherence, video quality, and faster speed. We validate this in our experiments in Table 3. Given $N$ tracks $\{(x_{t}^{n},y_{t}^{n})\}_{n=1}^{N}$ across $T$ temporal frames, the input track-conditioning signal $c_{m}\in\mathbb{R}^{T\times H/s\times W/s\times d}$ is constructed by placing visible track embeddings at spatially downsampled locations, where $s$ is the VAE spatial downsampling rate and $v[t,n]\in\{0,1\}$ indicates track visibility:
@@
Our lightweight track-encoding head performs $4\times$ temporal compression followed by a $1\times 1\times 1$ convolution. Prior methods adopt a ControlNet-style architecture (zhang2023adding; geng2025motion), which doubles FLOPs by duplicating network blocks. Instead, we directly concatenate the processed track embeddings with video latents, requiring only minor channel adjustments in the patchifying layer while leaving the core DiT architecture unchanged.
Training. We train the motion-guided teacher model through rectified flow matching objective (Liu2022FlowSA; lipman2022flow), where the forward process linearly interpolates between data $z_{0}$ and Gaussian noise $z_{1}\sim\mathcal{N}(0,I)$: $z_{t}=(1-t)z_{0}+tz_{1},\ t\in[0,1]$. The model is trained to predict the expected velocity fields with conditional flow matching loss $\mathcal{L}_{\text{FM}}$. One important limitation to note is that the model cannot inherently distinguish between occluded (non-visible) tracks and unspecified tracks, as both are represented by zero values. When a user releases controls during interaction, the model cannot determine whether the sudden zero values indicate occlusion or simply the absence of specification. This ambiguity occasionally leads to artifacts where objects abruptly appear or disappear. To address this issue, we introduce stochastic mid-frame masking with probability $p_{\text{mask}}=0.2$, where $c[t_{\text{rand}},:,:]=0$, for randomly selected mid-frame chunks $t_{\text{rand}}$. In practice, we first train the model without masking to establish strong track-following capability, and then fine-tune with stochastic masking to preserve coherence when track signals change intermittently.
@@
where $v_{\text{base}}=\alpha\cdot v(\varnothing,c_{m})+(1-\alpha)\cdot v(c_{t},\varnothing)$ and $\alpha=w_{t}/(w_{t}+w_{m})$ (we omit $z_{t}$ for brevity). We find that the joint guidance weights ($w_{t}=3.0$, $w_{m}=1.5$) provide a good balance: text conditioning enables realistic dynamics, even with sparse, flat-grid inputs, while track guidance preserves trajectories and maintains shape fidelity. We further analyze these effects in Sec. 4.3. Although this increases sampling cost from $2$ to $3$ function evaluations (NFE) per denoising step in the teacher model, our causal distillation (described in the next section) embeds all guidance into a single NFE, thereby eliminating this overhead in the student model.
@@ 3.2 Causal Distillation
Existing motion-controlled video diffusion models require around 50 denoising steps to generate high-quality video and also presume that the entire input motion trajectory is known before generation. In this section, we distill the slow teacher model into a causal video diffusion model, enabling real-time streaming of long videos with motion control. Our training pipeline starts from the Self Forcing paradigm huang2025self, but off-the-shelf, we find this exhibits large latency fluctuations due to varying attention window sizes and performs well only within the teacher’s training horizon (i.e., 81 frames), with quality quickly degrading when extrapolating to longer sequences. This section presents several key technical innovations to address these issues.
Attention sink in video diffusion model. Autoregressive generation with sliding-window attention, as employed in Self Forcing (huang2025self), is prone to quality degradation and drift during long-video extrapolation. To understand this failure, we visualize the self-attention maps in Figure 3 for both bidirectional and causal attention. Notably, many attention heads focus on the initial tokens corresponding to the input image. This phenomenon mirrors the observations in large language models xiao2023efficient, where initial tokens play a crucial role in stable streaming generation. Inspired by this, we adapt the attention sink concept to our video model by maintaining a local attention window while preserving the initial frame’s tokens as a fixed anchor during both training and inference. In our chunk-wise autoregressive setup, this sinking mechanism prevents the model from drifting during long-video extrapolation, even when trained only on short sequences limited by the teacher model.
Causal Adaptation. Following the initialization protocol from CausVid (yin2025causvid), we start our student model with the weights of the motion-guided teacher diffusion model and adapt it for causal attention architecture and few-step trajectory, using regression on ODE solution pairs sampled from the teacher. We train the model using attention mask with varying context window size and attention sink size for better generalization.
Self Forcing-Style Distillation. Following Self Forcing (huang2025self), we perform temporal autoregressive roll-out with distribution matching distribution (DMD) (yin2024one; yin2024improved). During roll-out, video chunks are generated sequentially, conditioned on previously self-generated outputs rather than ground truth using KV cache.
@@
In our approach, because the KV cache is continuously updated, RoPE values are assigned based on cache position rather than absolute temporal index. We apply KV cache rolling with attention sinks during both training and inference, implementing local attention without explicit attention masking and fully bridging the train–test gap, even in extrapolation scenarios. While TalkingMachines (low2025talkingmachines) also explores attention sinks, our method differs in two key aspects that enhance long-video stability. First, we eliminate the train-inference mismatch by explicitly simulating the extrapolation process during training, using self-rollout with a rolling KV cache and attention sinks. In contrast, TalkingMachines employs synchronized denoising with causal attention masks, which does not fully replicate the dynamics of autoregressive inference. Second, our training process ensures the teacher model always evaluates continuous video frames. The setup in TalkingMachines introduces a temporal discontinuity between the sink frame and subsequent frames, which can push the input outside the teacher’s pre-training distribution. By maintaining continuity, our teacher provides more robust scores for distillation.
@@
Intuitively, the driving gradient for the few-step causal generator $G_{\theta}$ comes from the difference between the estimated real and fake scores. To transfer the high-fidelity control of our joint guidance into the student without inference overhead, we define the target real score $s_{\text{real}}$ using the frozen teacher $f_{\phi}$ with joint guidance (omitting $z_{t}$ for brevity):
$s_{\text{real}}=s_{\text{base}}+w_{t}\cdot(f_{\phi}(c_{t},c_{m})-f_{\phi}(\emptyset,c_{m}))+w_{m}\cdot(f_{\phi}(c_{t},c_{m})-f_{\phi}(c_{t},\emptyset)),$ (4)
where $s_{\text{base}}$ follows the weighting defined in Eq. 2. In contrast, we parameterize the fake score $s_{\text{fake}}$ without using any CFG through a trainable critic $f_{\psi}$ (which approximates the generator’s score): $s_{\text{fake}}=f_{\psi}(c_{t},c_{m})$. This configuration effectively “bakes” the computational cost of the teacher’s multi-term guidance into the distillation objective, allowing the student generator to replicate the high-quality joint-guided distribution with a single function evaluation.
We update the generator ($G_{\theta}$) and fake score estimator ($f_{\psi}$) with a ratio of 1:5 (yin2024improved), allowing the critic to better approximate generated distributions. To manage GPU memory, we adopt Self Forcing’s gradient truncation strategy: randomly sampling denoising step $k$ from $[1,K]$ and using only that step’s output for self-rollout. This maintains computation graphs only for final denoised embeddings in the KV cache, reducing memory while preserving critical gradient flow.